im familar with writing conic sections in there standard equation but there are 2 questions that have me stumped in peticular.here is one of them.
y^2-8y-8x+64=0
and im suppose to write this in standard form

Answer by stanbon(75887)

y^2-8y-8x+64=0
and I'm supposed to write this in standard form
-----------------
Complete the square of the y terms, as follows:
y^2 - 8y + 16 = 8x -64 + 16
(y-4)^2 = 8x - 48
(y-4)^2 = 8(x-6)
--------------------
This form shows you the vertex is at (6,4) and the
focus of the parabola is 8/4 = 2 units above the vertex.
-------------------------------
Cheers,
Stan H.